The ages of 10 grooms at their first marriage are listed below. Find the midquartile. Round to the nearest tenth.

Start 1 By 10 Matrix 1st Row 1st Column 35.1 2nd Column 24.3 3rd Column 46.6 4st Column 41.6 5st Column 32.9 6st Column 26.8 7st Column 39.8 8st Column 21.5 9st Column 45.7 10st Column 33.9 EndMatrix35.1  24.3  46.6  41.6  32.9  26.8  39.8  21.5  45.7  33.9 
Answer #1

we arrange the data in ascending order

21.5 24.3 26.8 32.9 33.9 35.1 39.8 41.6 45.7 46.6  

^    ^ ^   

Q1 median Q3

we find out

Q1 and Q3

since n=10 ,

median is the mean of 5th and 6th observation

Q1 is the median of first 5 observation , that is third obs

Q1 = 26.8

Q3= 41.6

mid range = (Q1 +Q3)/2 = (26.8+41.6)/2 = 34.2